How many moles of carbon are in 2.5g of ethanol? 


1
Expert's answer
2022-05-02T04:46:03-0400

n(C2H5OH) = m(C2H5OH) / Mr(C2H5OH)

n(C2H5OH) = 2.5 g / 46.07 g/mol = 0.05 moles


1 mol of C2H5OH contains 2 mol of atomic carbon

n(C) = 2n(C2H5OH)

n(C) = 2 * 0.05 moles = 0.1 moles